AC-LGAD working group meeting

US/Eastern
Description

Meeting for all AC-LGAD subsystems to discuss topics of common interest (e.g. sensors, ASICs, testbeams).

 

This meeting will use the TOF Zoom link: https://bnl.zoomgov.com/j/1617546118?pwd=qNzxLqF8Q4Mj3RerAZdVSELzEgEQzV.1

 

General agenda:

1) Short updates from institutions on AC-LGAD efforts

2) Discussion of plan/goals for potential SPS testbeam et al.

3) Question which need to be addressed before a new sensor production can be ordered.

Attendees: Alex, Satoshi, Ashik, Elke, Grigory, Jennifer, John, Masashi, Mathieu, Olga, Prithwish, Takashi, Taku, Tommy, Yu, Zhangbu, Beni, Gabriele, Dominique, Kentaro

Update on SPS testbeam: RE47 (our activity number) has a week in the "draft" schedule in mid-July, but I have NOT received any official email from them confirming this, so it is still in "draft" form on the schedule. I have already reached out to Martin for an update.

 

Satoshi's Update:

  • DESY testbeam analysis updates
    • 2 digitizers -> 3 LGADs to use for timing reference.
    • Data successfully calibrated and initial amplitudes look reasonable for each channel.
    • Position reconstruction data shows anomalous behavior near strip for two different reconstruction methods --> still under study.
    • Hit position correlation between sensors (with first sensor as reference) shows some deviation from slight misalignment of DUTs.
    • Timing difference data between hits in different seniors show very strange two-band structure - needs further study to understand.
  • KEK testbeam (parasitic) just started with pixel sensors (20um and 30um) and strips (50um) - 6 GeV beam. (now until February. 25th)
  • Tohoku-RARiS, electrons at 800 MeV happening March 3rd to 6th -- goal is to look at other strip sensors geometries and follow-up on questions from previous campaigns.
    • "plan B" sensors. --> looking at shorter strips with larger pitch, degrade position resolution in favor of timing resolution, since the latter drives the TOF requirement.
  • KEK 6.5 GeV (main) beam planned for late Fall (Oct. to Dec.) with a primary goal of testing full chain with strips + FCFDv1.2
  • Elke: Question to "plan B" sensors --> what about impact on FCFD design if this route needs to be pursued? --> Would require re-design of FCFD analog front-end, so this option needs much more discussion. --> Goal now is just to study the different sensor and see what is gained with other option. --> different length contributes to different capacitance, and may also reduce active area.
  • Beni: Half strip length with "plan B" means double the # of channels? --> No, pitch is increased, so it cancels out.
  • Zhangbu: do we have any study of dead area as a function of different pitch? --> not at the moment.

 

Taku's Update:

  • Focus on updates on investment from Japanese funds (see slides for detailed breakdown).
    • Sensor productions, lab equipment (bonding, clean room, digitizers, etc.).

 

Dominique's update:

  • EICROC1 received at OMEGA, and initial testing work is underway.
  • IR laser data analysis underway at IJCLab.
  • Experiencing issues with new firmware for EICROC0.

 

Grigory's update:

  • UIC can characterize sensors, perform TCT, etc. --> participated in previous testbeam campaigns.
    • Will be important as we receive new productions in the future.
  • Have cold blocks for cooling DUTs --> can be used in testbeam to provide stable temps for DUTs.
  • UIC wants to be involved in helping with databasing for sensors --> Simone already has a table, but maintainence will be important given the number of sensors we have on-hand, and how many will come in the future.
  • A storage spot has been put together at BNL on GPFS02 to help with storage of teatbeam data.
    • We will need to look at how to manage this longterm as we don't have unlimited storage (~10TB per campaign).

 

Jenni's update:

  • Contributions aimed at TOF, with a focus on fTOF. 
  • Question about hybridization and QA --> all done at Oak Ridge? Would be good to have multiple locations.
    • Jenni: have IR laser, probe station, analyzer, and small clean room. --> could have a setup at Hawaii to provide additional QA and testing for production.
  • Mathieu: ASIC probing can be done at Oak Ridge, equipment exists already.
  • Alex: Bump-bonding planned at ORNL, or at a vendor. --> Vendor.

 

BNL updates:

  • Alex --> updates on EICROC0 + AC-LGAD testing. --> no questions
  • Prithwish --> updates on RBv1 setup and power board. 
    • 3 power boards made, 2 tested.
    • RBv1 setup is up at BNL, waiting FPGA programming from Tonko, but the components are ready.
    • Mathieu: Why not use a FELIX? --> not ready yet, the current card is a light version of this.
  • Gabriele Giacomini --> AC-LGAD productions at BNL 
    • Some wafers almost ready - just needs passivation layer. --> 4x4 sensors, and some strip production (already being distributed).
    • Elke: BNL setup was used to check various things we couldn't do at HPK. --> addressing double-metal, and different technologies, pure AC-LGAD vs. some other approaches (implant embedded in resistive layer to reduce cross-talk). --> Some summary update is needed to come to closure point on R&D.
    • Elke: any news from Taiwan? --> need to meet to finalize process, and then prepare first batch for production --> no timeline yet.

 

LBL updates from Yu Hu:

  • update on JLAB testbeam results.
  • Elke: This poor timing resolution is due to the incorrect bias voltage? --> Yes, and the DESY result already show a larger signal amplitude, so we will likely see better timing resolution from that campaign to confirm.

 

 

 

 

 

 

 

 

 

There are minutes attached to this event. Show them.